# Zeno Effect Suppression of Gauge Drift in Quantum Simulations

### v2

@@ I Introduction
The method presented in this paper concerns digital quantum simulation. Ideally, any method to simulate a non-Abelian gauge theory of interest, such as quantum chromodynamics (QCD), would only simulate the gauge invariant (i.e. physical) states. Methods of this kind have been proposed [46, 47], which encode only physical degrees of freedom. While these methods don’t have to deal with unphysical states, they are fundamentally more challenging. Thus, many methods simulate all states in the Hilbert space of the system, and must contend with keeping the system within the exponentially smaller physical subspace. Crucially, work has been done to develop methods of time evolution that respect gauge invariance [13, 48]. Thus, ideally time evolving an initial state that is constructed to be physical subspace would keep the system within the physical subspace. However, quantum noise, gate errors, and various errors from approximations can lead to the system picking up contributions from unphysical states. Previous work has addressed this problem in a variety of ways. Perhaps the most straightforward is to simply accept the unphysical contributions on the premise that these contributions are reduced as the time step is decreased [15]. Another method is dynamical decoupling [49], a concept from quantum control theory where local symmetries are maintained via periodic driving. Also, oracles have been developed [50] to check for violations of Gauss’ law for Abelian lattice gauge theories, with similar oracles for non-Abelian theories under development.
@@ II.1 Projection via the Rodeo Algorithm
To make clear how the rodeo algorithm works, we will write out how it works for one ancillary qubit. Starting with the system in the state |$\ket{\psi_{I}}$ and the ancillary qubit initialized to |$\ket{1}$, we have
@@
A successful implementation of the rodeo algorithm measures every ancillary qubit in the state |$\ket{1}$ at the end. In the single ancillary qubit example above, this results in the transformation |$\ket{\psi_{I}}\rightarrow\frac{1}{2}(I+e^{-i(H_{obj}-E)t})\ket{\psi_{i}}$ where $E$ is the energy of the target eigenstate and $H_{obj}$ is the Hamiltonian of the system. Decomposing our initial state into energy eigenstates |$\ket{\psi_{i}}=\sum\limits_{n}c_{n}\ket{n}$, we can see that the probability for each energy eigenstate $n$ transformations from $|c_{n}|^{2}$ before Rodeo projection to $|c_{n}|^{2}|\frac{1}{2}(1+e^{-i(E_{n}-E)t})|^{2}=|c_{n}|^{2}\cos^{2}((E_{obj}-E)\frac{t}{2})$ afterwards. Thus, the probability for the system to be in a particular energy eigenstate $E_{obj}$ is multiplied by a factor $\cos^{2}((E_{obj}-E)\frac{t}{2})$, which is notably 1 if $E=E_{obj}$ and less than 1 otherwise. This is how the rodeo algorithm works: it constructs an approximate version of the state of energy $E$ by suppressing all other states with a $\cos^{2}(\frac{1}{2}t\Delta E)$ factor. Stronger suppression of contributions from unwanted energy eigenstates can be accomplished with multiple rounds of this algorithm, as discussed in [53, 54, 55]. Note that in order to implement a round of the rodeo algorithm, a time $t$ must be chosen. If the times are chosen strategically for the various rounds, remarkably strong suppression can be achieved with very modest resources [56].
@@
We note here that the Rodeo projection will fail to project onto the state of interest with finite probability (i.e. if any ancillary qubit is not measured in |$\ket{1}$), where failure means that the resulting state is orthogonal to the target state. Crucially, we note that the probability of a successful projection is proportional to the overlap of the initial state of the system with the target state of the projection. As we plan to use this projection to leverage the Zeno effect, which means making sufficiently many measurements so that the overlap with the target subspace is close to unity, the rate of failure can be made negligibly small. The chance of failure serves as the probabilistically defined time span for which our method works; for example, if the probability for a failed projection is one in a million, we would expect one failure within one million successive projections.
The rodeo algorithm is designed to construct eigenstates of a system’s Hamiltonian. For our purposes, we use the ideas of this algorithm for a different hermitian operator, that we will call $G^{2}$:
@@
From the above definition of the Gauss’ law operators, we see that the physical states are eigenstates of $G^{2}$ with eigenvalue 0. Furthermore, all the other eigenstates will be unphysical and with eigenvalues greater than 0, since $G^{2}$ is positive definite by construction. It is this spectrum of $G^{2}$ that we will exploit when using the rodeo algorithm: instead of constructing some eigenstate of a given Hamiltonian, we will construct a physical state by constructing a ground state of $G^{2}$ via the Rodeo algorithm. This means replacing $H_{obj}$ in 1 with $G^{2}$ as well as removing the phase rotation gates (since we would be aiming for an eigenstate with an eigenvalue of 0, so the required phase rotation is 0*t = 0).
@@ III Performance
To study this we consider a small “toy” model. It is a pure $\mathbb{Z}_{2}$ gauge theory in 2 spatial dimensions [59]. We simulate this on a 2x2 lattice with periodic boundary conditions. The purpose of using such a small simulation is simply that it allows one to explore the method with very little cost. The Hamiltonian for the toy model is
$H=-k\sum\limits_{j}\square_{j}-g\sum\limits_{i}\sigma^{x}_{i}$ (7)
@@
where $\square_{j}$ is the $j$th plaquette operator, $\sigma^{x}_{i}$ is the Pauli x operator on the $i$th link, and $k$ and $g$ are parameters. A plaquette operator in this context is a product of the $\sigma^{z}$ operators living on the four sides of a plaquette in the lattice.
In this system, the Gauss’ law operators are constructed by taking the product of the $\sigma^{x}$ operators living on the four links emanating from a site on the lattice. We note here that while our method in general works by constructing an operator $G^{2}$ consisting of a sum of the square of all Gauss’ law operators, we have to modify this approach for this toy model as the square of a product of $\sigma^{x}$ operators from different links is manifestly the unity operator. Instead, we use a direct sum of the four Gauss’ law operators of this system, which we will call $G_{tot}$. We also note that for this system, we define a physical state as one satisfying
$G_{a}\ket{\psi}=\ket{\psi}$ (8)
for $a=1,2,3,4$ denoting the associated site. As each $G_{a}$ has a 2-level spectrum (1 and -1), we see that $G_{tot}$ has a 3-level spectrum (-4, 0, and 4), with the physical states satisfying $G_{tot}\ket{\psi}=4\ket{\psi}$.
@@
Our aim is to study gauge drift, so we include in our toy model an artificial gauge drift operator $D$ with each time step. We base our constructed gauge drift operators off of eq. (6); specifically, we connect each of the, in our case 32, physical states of the system to a randomly selected unphysical state in exactly the way laid out in eq. (6). All other states are unchanged. This provides a way for the system to drift out of the physical subspace of its Hilbert space in a unitary way, and a parameter, $\epsilon$, to control the strength of this drift. For our simulations, we set this parameter to $\epsilon=0.01$. Note that each time our method calls for time evolution, we select a different set of 32 unphysical states to appear in the gauge drift operator so as to not bias any unphysical states.
@@
With this defined, the algorithm can be used for the toy model. Each time step takes a state |$\ket{\psi}$ to $G_{i}DU_{t}\ket{\psi}$ where $G_{i}$ is a randomly chosen gauge transformation, $D$ is a gauge drift operator, and $U_{t}=\exp{-iHt}$ ) is the standard time evolution operator. The remaining part of the algorithm is the rodeo projection, which occurs after a set number of time steps (determined in the calibration phase). As discussed above, a successful projection takes a state |$\ket{\psi}$ to $\frac{1}{2}(1+\exp(-i(G_{tot}-4)t_{r}))\ket{\psi}$ where here we note that the time used within the projection $t_{r}$ is in general different from the time step size $t$. Recall that all physical states are eigenstates of $G_{tot}$ with an eigenvalue of 4, so they are unaffected while unphysical states will suffer destructive interference.
@@
Figure 2: Comparing gauge drift suppression methods, including no suppression methods (green), a random gauge transformation after each time step (blue), and a random gauge transformation followed by a rodeo projection after each time step (orange). The right plot is the same as the left plot without the green plot so as to zoom in on the behavior of the other two.
@@
Figure 3: Comparing performance of regimes with no gauge transformations and various frequencies of rodeo projections with regimes with gauge transformations and various frequencies of rodeo projections
@@ IV Implementation
It is important to consider the implementation of this method, as well as their costs. We will start with a conisderation of the implementation of the gauge transformations, which is discussed in [58]. First of all, there is the cost of sampling the gauge transformations, which includes a small number of ancillary qubits; notably, this cost does not scale with the size of the system [48, 58]. When it comes to implementing the sampling, we make use of a common method of sampling uniformly from the Haar measure of a group $G$ wherein a small number (e.g. 2) elements of $G$, call them $g_{1}$ and $g_{2}$, are chosen and then strings (of a chosen length) of these elements are randomly sampled, e.g. $g_{1}g_{2}g_{2}g_{1}g_{2}$. Importantly, this method of sampling elements of $G$ converges to the Haar measure in the limit of long strings. For the present purpose, however, we note that we do not rely on a truly uniform sampling of gauge transformations; thus, relatively short strings can be used, in a similar way to the circuit described in [60].
Next, we consider the implementation of the gauge transformations, which, as the original authors of this method noted in ref.  [58], ”requires about as many G-multiplication gates as the original quantum simulation did”. Thus, the total cost of using gauge transformations in this method is about double what it would be for a straightforward simulation.
@@
Additionally, the implementation and cost of the gauge projections needs to be considered. We have suggested that the rodeo projection method appears to be useful. Note that previous work has shown that the rodeo algorithm can be optimized and made far more efficient [56] via considered choices for the parameter $t_{r}$. We suggest using this optimized approach makes sense for our purposes as well. This work notes that the dominant cost of the rodeo algorithm is the controlled time evolution gate (in our case it is then the controlled U gate with $U=\exp(-iG^{2}t_{r})$); thus, they argue that the total time to conduct a rodeo projection acts as a proxy for computational costs and then proceeds to lay out a method for selecting the $t_{r}$s that maximize efficiency of the algorithm given fixed computational resources [56]. We will also note that constructing $G^{2}$ will be much more difficult than $H$ so our version of a rodeo projection suffers a much heftier cost in construction of the required gates, but so will alternative projection schemes.
@@
Furthermore, the projections themselves require a number of ancillary qubits. Notably, the projections only use one ancillary qubit at a time, so if mid-circuit measurements are allowed then only one ancillary qubit is required, as it can be recycled again and again. If this is not the case, then a significant number of ancillary qubits must be used. Specifically, we estimate one projection would need as many ancillary qubits as sites in the lattice.

### v4

@@ I Introduction
The method presented in this paper concerns digital quantum simulation. Ideally, any method to simulate a non-Abelian gauge theory of interest, such as quantum chromodynamics (QCD), would only simulate the gauge invariant (i.e. physical) states. Methods of this kind have been proposed which solve the Gauss law, either eliminating the gauge fields [46, 47, 48] or the matter fields [49, 50, 51, 52, 53]. Another method uses dual formulations of the gauge theories to enforce the Gauss law [54, 55, 56]. While the methods discussed above don’t have to deal with unphysical states, their physical implementation may be more challenging.
Thus, many methods simulate all states in the Hilbert space of the system, and must contend with keeping the system within the exponentially smaller physical subspace. Crucially, work has been done to develop methods of time evolution that respect gauge invariance [13, 57]. Thus, ideally time evolving an initial state that is constructed to be physical would keep the system within the physical subspace. However, quantum noise, gate errors, and various errors from approximations can lead to the system picking up contributions from unphysical states. Previous work has addressed this problem in a variety of ways. Perhaps the most straightforward is to simply accept the unphysical contributions on the premise that these contributions are reduced as the time step is decreased [15]. Another method is dynamical decoupling [58], a concept from quantum control theory where local symmetries are maintained via periodic driving. Also, oracles have been developed [59] to check for violations of Gauss’ law for Abelian lattice gauge theories, with similar oracles for non-Abelian theories under development.
@@ II.1 Projection via the Rodeo Algorithm
The rodeo algorithm aims to construct, to good approximation, an eigenstate of a system with a chosen energy by ”shaking off” amplitudes of other, unwanted eigenstates, like a bull shaking off its rider at a rodeo. The algorithm accomplishes this task with ancillary qubits as shown in the circuit diagram of Fig. 1. To see how the rodeo algorithm constructs the desired eigenstate, we will write out how it works for one ancillary qubit. Starting with the system in the initial state |$\ket{\psi_{I}}$ and the ancillary qubit initialized to |$\ket{1}$, we have
@@
$\begin{bmatrix}\frac{I}{\sqrt{2}}&\frac{I}{\sqrt{2}}\\ \frac{I}{\sqrt{2}}&-\frac{I}{\sqrt{2}}\end{bmatrix}\begin{bmatrix}I&0\\ 0&Ie^{iEt}\end{bmatrix}\begin{bmatrix}I&0\\ 0&e^{-iH_{obj}t}\end{bmatrix}\begin{bmatrix}\frac{1}{\sqrt{2}}\ket{\psi_{I}}\\ -\frac{1}{\sqrt{2}}\ket{\psi_{I}}\end{bmatrix}=$ (3)
$\begin{bmatrix}\frac{I}{\sqrt{2}}&\frac{I}{\sqrt{2}}\\ \frac{I}{\sqrt{2}}&-\frac{I}{\sqrt{2}}\end{bmatrix}\begin{bmatrix}\frac{1}{\sqrt{2}}\ket{\psi_{I}}\\ -\frac{1}{\sqrt{2}}e^{-i(H_{obj}-E)t}\ket{\psi_{I}}\end{bmatrix}=$ (4)
@@
where above, $H_{obj}$ is the Hamiltonian of the system of interest (ie the object system) and $E$ is the desired energy of the constructed state. Equation (2) writes out the actions of (reading from right to left) the Hadamard, controlled time evolution, phase shift, and Hadamard gates, respectively, on the initial state. Note that the first Hadamard gate gives equal amplitude, up to a sign, to the up (ie |$\ket{0}$) and down (ie |$\ket{1}$) components of the ancillary qubit (equation (3)), then the controlled time evolution and phase shift gates give a phase only to the down component (equation (4)), and finally the last Hadamard gate causes interference between the up and down components (equation (5)).
The final step in the rodeo algorithm is to measure the ancillary qubit(s). A successful implementation of the rodeo algorithm measures every ancillary qubit in the state |$\ket{1}$ at the end. In the single ancillary qubit example above, this results in the transformation
|$\ket{\psi_{I}}\rightarrow\frac{1}{2}(I+e^{-i(H_{obj}-E)t})\ket{\psi_{I}}$ (6)
To see the benefits of this transformation, we decompose our initial state into energy eigenstates |$\ket{\psi_{I}}=\sum\limits_{n}c_{n}\ket{n}$ where $H_{obj}\ket{n}=E_{n}\ket{n}$. The amplitude of each energy eigenstate then transforms as
$c_{n}\rightarrow\frac{1}{2}c_{n}(1+e^{-i(E_{n}-E)t})$ (7)
which means the probability of each energy eigenstate transforms as
$|c_{n}|^{2}$ $\rightarrow|c_{n}|^{2}|\frac{1}{2}(1+e^{-i(E_{n}-E)t})|^{2}$ (8)
$\rightarrow|c_{n}|^{2}\cos^{2}((E_{n}-E)\frac{t}{2})$ (9)
@@
Thus, the probability for the system to be in a particular energy eigenstate $E_{n}$ is multiplied by a factor $\cos^{2}((E_{n}-E)\frac{t}{2})$, which is notably 1 if $E=E_{n}$ and less than 1 otherwise. This is how the rodeo algorithm works: it constructs an approximate version of the state of energy $E$ by suppressing all other states with a $\cos^{2}(\frac{1}{2}t\Delta E)$ factor. Stronger suppression of contributions from unwanted energy eigenstates can be accomplished with multiple rounds of this algorithm, as discussed in [62, 63, 64]. Note that in order to implement a round of the rodeo algorithm, a time $t$ must be chosen. If the times are chosen strategically for the various rounds, remarkably strong suppression can be achieved with very modest resources [65].
@@
We note here that the rodeo algorithm will fail to project onto the state of interest with finite probability (i.e. if any ancillary qubit is not measured in |$\ket{1}$), where failure means that the resulting state is orthogonal to the target state. Crucially, we note that the probability of a successful projection is proportional to the overlap of the initial state of the system with the target state of the projection. As we plan to use this projection to leverage the Zeno effect, which means making sufficiently many measurements so that the overlap with the target subspace is close to unity, the rate of failure can be made negligibly small. The chance of failure serves as the probabilistically defined time span for which our method works; for example, if the probability for a failed projection is one in a million, we would expect one failure within one million successive projections.
So far, we have focused on the rodeo algorithm due to its key feature of leaving desired amplitudes (ie the amplitude of the desired energy eigenstate) untouched while suppressing all other amplitudes. Here, we note that there is no reason this algorithm works only for the construction of the eigenstates of the Hamiltonian. Thus, going forward, we will refer to rodeo projections as any projection of the form expressed in Fig. 1 with some hermitian operator in place of the Hamiltonian.
Recall that our goal is to use frequent projections onto the physical subspace to maintain gauge invariance of the system via the Zeno effect. With this in mind, we now discuss the form of a rodeo projection that we will use to project onto the physical subspace. We replace the Hamiltonian with a new hermitian operator that we will call $G^{2}$:
@@
From the above definition of the Gauss’ law operators, we see that all physical states are eigenstates of $G^{2}$ with eigenvalue 0. Furthermore, all the other eigenstates will be unphysical and with eigenvalues greater than 0, since $G^{2}$ is positive definite by construction. It is this spectrum of $G^{2}$ that we will exploit when using rodeo projections: instead of constructing some eigenstate of a given Hamiltonian, we will construct a physical state by constructing a ground state of $G^{2}$ via rodeo projections. This means replacing $H_{obj}$ in Fig. 1 with $G^{2}$; furthermore, as the eigenvalue of the desired eigenstate (ie a physical state) is 0, we will remove the phase shift gates, since the required phase shift is $00*t=$ 0.
This is the core of the method outlined in this paper: a rodeo projection with this $G^{2}$ operator will leave all physical amplitudes untouched while suppressing all unphysical amplitudes. Doing this projection frequently enough will keep the system in a close approximation of a gauge invariant state via the Zeno effect.
@@ III Performance
Figure 2: The diagram of the toy model with 4 sites and 8 links
@@
To study our algorithm, we consider a small “toy” model. It is a pure $\mathbb{Z}_{2}$ gauge theory in 2 spatial dimensions [68]. We simulate this on a 2x2 lattice with periodic boundary conditions (see Figure 2). The purpose of using such a small simulation is simply that it allows one to explore the method with very little cost. The Hamiltonian for the toy model is
$H=-g\sum\limits_{i=A}^{H}\sigma^{x}_{i}-k\sum\limits_{j=i}^{4}\square_{j}$ (13)
@@
where $g$ and $k$ are parameters, $\sigma^{x}_{i}$ is the Pauli x operator on the $i$th link, and $\square_{j}$ is the $j$th plaquette operator (where the $j$th plaquette is defined as the plaquette with the $i$th site in the bottom left corner). A plaquette operator in this context is a product of the $\sigma^{z}$ operators living on the four sides of a plaquette in the lattice; for example, the first plaquette operator is
$\square_{1}=\sigma^{z}_{A}\sigma^{z}_{B}\sigma^{z}_{C}\sigma^{z}_{G}$ (14)
In this system, each site has an associated local gauge transformation $g_{i}$, constructed by taking the product of the $\sigma^{x}$ operators living on the four links emanating from the $i$th site on the lattice:
$g_{1}$ $=\sigma^{x}_{A}\sigma^{x}_{C}\sigma^{x}_{D}\sigma^{x}_{E}$ (15)
$g_{2}$ $=\sigma^{x}_{B}\sigma^{x}_{C}\sigma^{x}_{D}\sigma^{x}_{F}$ (16)
$g_{3}$ $=\sigma^{x}_{A}\sigma^{x}_{E}\sigma^{x}_{G}\sigma^{x}_{H}$ (17)
$g_{4}$ $=\sigma^{x}_{B}\sigma^{x}_{F}\sigma^{x}_{G}\sigma^{x}_{H}$ (18)
We note here that our gauge group $\mathbb{Z}_{2}$ is discrete, which means we cannot formulate Gauss’ law via generators of gauge transformations. Instead, we note that the Hilbert space of the system is split into sectors defined by the eigenvalues of the gauge transformations in equations (15)-(18). Note that each $g_{i}$ has eigenvalues ±$\pm 1$. With this, we define a physical state as a state |$\ket{\psi}$ that is unchanged by all local gauge transformations:
$g_{i}\ket{\psi}=\ket{\psi}\;\forall i$ (19)
At this point we will note that in previous sections we described our algorithm using a $G^{2}$ operator (defined in equation (10)), but this was for systems governed by compact Lie gauge groups. Here, we have the discrete group $\mathbb{Z}_{2}$, which requires two small alterations. Firstly, we replace the $G^{2}$ operator with a new operator $G_{tot}$:
$G_{tot}=g_{1}+g_{2}+g_{3}+g_{4}$ (20)
We note that the spectrum of $G_{tot}$ is $-4$,00, and $4$; furthermore, using equations (19) and (20), we have that
$G_{tot}\ket{\psi}=4\ket{\psi}$ (21)
for all physical states. This $G_{tot}$ operator serves the same purpose as the $G^{2}$ operator: every physical state is an eigenstate with the same eigenvalue (4 for $G_{tot}$ and 0 for $G^{2}$) and all unphysical states have different eigenvalues. Thus, $G_{tot}$ will function the same in our algorithm for our discrete group as $G^{2}$ will for compact Lie gauge groups.
Here, we must highlight the second alteration needed for our discrete group. Figure 1 shows phase rotation gates as part of the circuit diagram, but we argued that they are not required in our algorithm in section IIA due to physical states having an eigenvalue of 0. As just discussed, this is not true for our toy model, so we must include the phase rotation gates with $E=4$, the eigenvalue of physical states for our system.
@@
Now, our aim is to study gauge drift, so we include in our toy model an artificial gauge drift operator $D$ with each time step. We base our constructed gauge drift operators off of eq. (12); specifically, we connect each of the, in our case 32, physical states of the system to a randomly selected unphysical state in exactly the way laid out in eq. (12). All other states are unchanged. This provides a way for the system to drift out of the physical subspace of its Hilbert space in a unitary way, and a parameter, $\epsilon$, to control the strength of this drift. For our simulations, we set this parameter to $\epsilon=0.01$. Note that each time our method calls for time evolution, we select a different set of 32 unphysical states to appear in the gauge drift operator so as to not bias any unphysical states.
@@
With this defined, the algorithm can be used for the toy model. Each time step takes a state |$\ket{\psi}$ to $\tilde{g}_{i}DU_{t}\ket{\psi}$ where $\tilde{g}_{i}$ is a randomly chosen gauge transformation, $D$ is a gauge drift operator, and $U_{t}=\exp{-iHt}$ ) is the standard time evolution operator. Note that $\tilde{g}_{i}$ represents any gauge transformation on the system, including the $g_{i}$’s in equations (15)-(18) as well as any product of them. The remaining part of the algorithm is the rodeo projection, which occurs after a set number of time steps (determined in the calibration phase). As discussed above, a successful projection takes a state |$\ket{\psi}$ to $\frac{1}{2}(1+\exp(-i(G_{tot}-4)t_{r}))\ket{\psi}$ where here we note that the time used within the projection $t_{r}$ is in general different from the time step size $t$. Recall that all physical states are eigenstates of $G_{tot}$ with an eigenvalue of 4, so they are unaffected while unphysical states will suffer destructive interference.
@@
Figure 3: Comparing gauge drift suppression methods, including no suppression methods (dash-dot line), a random gauge transformation after each time step (dashed line), and a random gauge transformation followed by a projection after each time step (solid line). The right plot is the same as the left plot without the dash-dot plot so as to zoom in on the behavior of the other two.
@@
Figure 4: Comparing performance of regimes with no gauge transformations and various frequencies of projections (circle markers) with regimes with gauge transformations and various frequencies of projections (square markers)
@@ IV Implementation
It is important to consider the implementation of this method, as well as the associated costs. As our method involves two main operations (rodeo projections and gauge transformations), we will take them in turn.
IV.1 Implementing Rodeo Projections
With a circuit diagram of Fig. 1 plus the small alterations discussed in section IIA in hand, there remain two main questions when it comes to the implementation of our version of rodeo projections: how should the time parameters, call them $t_{r}$, be chosen, and how many ancillary qubits per projection are required?
@@ IV.1 Implementing Rodeo Projections
Firstly, we address the issue of the selection of the $t_{r}$ parameters. Here, we note that previous work has shown that the rodeo algorithm can be optimized and made far more efficient [65] via considered choices for the parameter $t_{r}$. We suggest that this optimized approach makes sense for our purposes as well. This work notes that the dominant cost of the rodeo algorithm is the controlled time evolution gate. In our case it is then the controlled U gate with $U=\exp(-iG^{2}t_{r})$. We will also note that constructing $G^{2}$ will be much more difficult than $H$ so our version of a rodeo projection suffers a much heftier cost in construction of the required gates, but so will alternative projection schemes. Due to the controlled U gate dominating costs, the authors of [65] argue that the total time to conduct a rodeo projection acts as a proxy for computational costs and then proceeds to lay out a method for selecting the $t_{r}$s that maximizes efficiency of the algorithm given fixed computational resources [65]. The authors note that choosing the $t_{r}$s randomly (eg via Gaussian distributions) can lead to exponentially large fluctuations in the suppression of undesired amplitudes; furthermore, they show that considered choices of the $t_{r}$s, specifically choosing $t_{r}$s that vary across exponentially many timescales, can eliminate the problem of these large fluctuations, thus leading to a more efficient implementation of the rodeo projection.
@@
Additionally, we must address the question of qubit costs. Notably, if mid-circuit measurements are allowed, then only one ancillary qubit is required, as it can be reused again and again. If this is not the case, then we estimate one projection would need roughly as many ancillary qubits as sites in the lattice. Future work aims to study how the qubit per projection requirement scales with system size.
IV.2 Implementing Gauge Transformations
It is important to discuss the practical implementation of random gauge transformations as [67] outlines in section III of their paper. First, we must sample from the Lie group $G$ of gauge transformations. One way is to construct an ancillary ’clock’ register that increases its value by 1 after each gauge transformation is conducted in the course of our algorithm. Then a random circuit (eg of the kind developed in [69]) will take the value of the clock and select a gauge transformation for each site and store it on an ancillary G-register. Once sampled, these gauge transformations can be performed on the system.
The first major cost consideration of this process is qubit costs. The sampling process requires two additional registers (the clock and the G-registers); notably, this qubit cost does not scale with the size of the system. Furthermore, the actual implementation of the gauge transformations on the system ”requires about as many G-multiplication gates as the original quantum simulation did” [67].
A second cost consideration is the depth of the random circuit. Intuitively, shorter circuits will result in less fair sampling from the Haaar measure of the Lie group than longer circuits. Luckily, we are helped significantly by the fact that we have no strict requirements on fair sampling from the Haar measure. Here, we point to a common method of sampling uniformly from the Haar measure of a group $G$, which is to select a small number (eg 2) of elements of $G$, call them $g_{1}$ and $g_{2}$, and construct strings of these elements of a chosen length (an example of a string of length 5 is $g_{1}g_{2}g_{2}g_{1}g_{2}$). Importantly, this method of sampling elements of $G$ converges to the Haar measure in the limit of long strings. This tells us that unfair sampling can be made more fair by conducting multiple unfairly sampled gauge transformations. Thus, we can use a short random circuit when sampling gauge transformations and then perform multiple gauge transformations in a row if need be to increase the fairness of the sampling. Future work aims to clarify the relationship between the fairness of sampling and the performance of the algorithm.